OpenStudy (anonymous):

Let a,b,c non-negativenumber such that a+b+c=3. prove that: 1/(2ab^2 +1) + 1/(2bc^2 +1) + 1/(2ca^2 +1) >= 1

OpenStudy (anonymous):

what u means mukushla??

OpenStudy (anonymous):

i mean i'll post my answer later i gotta go now...

OpenStudy (anonymous):

i dont wanna make a complete solution try to work it out 1. \[ \frac{1}{2ab^{2}+1}+\frac{1}{2bc^{2}+1}+\frac{1}{2ca^{2}+1}\geq 1\:\Longleftrightarrow\:\frac{ab^{2}}{2ab^{2}+1}+\frac{bc^{2}}{2bc^{2}+1}+\frac{ca^{2}}{2ca^{2}+1}\leq 1 \] (why?) 2. By AM-GM we have \(2ab^{2}+1\geq 3\sqrt[3]{a^{2}b^{4}}\) hence inequality becomes to \[ \frac{ab^{2}}{2ab^{2}+1}+\frac{bc^{2}}{2bc^{2}+1}+\frac{ca^{2}}{2ca^{2}+1}\leq\frac{1}{3}(\sqrt[3]{ab^{2}}+\sqrt[3]{bc^{2}}+\sqrt[3]{ca^{2}}) \] (why?) 3.and Using Chebyshev gives \[ \frac{1}{3}(\sqrt[3]{ab^{2}}+\sqrt[3]{bc^{2}}+\sqrt[3]{ca^{2}})\leq\frac{a+b+c}{3}= 1 \] (why?)

OpenStudy (anonymous):

U can study about Chebyshev's inequality and some other inequalities from this usefull resource

OpenStudy (anonymous):

OK, million thanks @mukushla...

OpenStudy (anonymous):

yw :)